Kohl Company makes computer bags. Its sales budget for four months is:

Sales
                     Month
  

                      December 15,000

                      January 20,000

                       February 40,000

                      March 60,000

Kohl's policy is that ending inventory of finished bags should equal 30% of the next
month's sales. Beginning inventory (December 1) is 4,500 bags. Budgeted production cost is $5.00

Each bag required 1.5 yards of ballistic nylon. The ending inventory policy for nylon is
that 20% of the following month's production needs must be on hand. On December 1, Kohl
had 3,300 yards of nylon in inventory. Budget Ballistic Nylon cost is $0.75/yard.

1. Compute the budgeted Ending Finished Goods Inventory for January:

2. Compute the Budgeted Beginning Finished Goods Inventory for January:

3. What is the budgeted COGM for bags in January?

4. Compute the budgeted Ending Materials Inventory for December:

5. What is the budgeted Cost of Goods Sold for January?
Answer #1
Dec Jan Feb Mar
Budgeted Sales 15000 20000 40000 60000
Add: Desired ending Inventory 6000 12000 18000
Total required 21000 32000 58000 60000
Less: beginning Inventory 4500 6000 12000
Units to be produced 16500 26000 46000
Raw material per unit 1.5 1.5 1.5
Yards required for production 24750 39000 69000
Add: Desired ending Inventory 7800 13800
Total required 32550 52800
Less: beginning Inventory 3300 7800
Required Purchases 29250 45000
Cost per yard 0.75 0.75
Budgeted cost 21937.5 33750
Hence, the answers are
Budgeted ending FG inventory = 12000 units
2.Beginning = 6000 units
3.Budgeted COGM in January = 26000*5 = $130,000
4.Ending material inventory = 7800 yards
5.COGS = 20,000*5 = $100,000